OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (anonymous):

circle equation formula is x^2+y^2=r^2

OpenStudy (anonymous):

(x-5)^2+(y-6)^2=9

OpenStudy (anonymous):

(x-2)^2+(y-6)^2=3^2

OpenStudy (anonymous):

exactly

OpenStudy (mertsj):

@ cnsu90 @MathEagle It is against code of conduct to post answers without attempting to help the asker learn how to do the problem

OpenStudy (anonymous):

sorry. i didnt know. :(